# Smart Glasses for CVI: Co-Designing Extended Reality Solutions to Support Environmental Perception by People with Cerebral Visual Impairment

### v1

@@ Abstract.
copyright: noneconference: The 27th International ACM SIGACCESS Conference on Computers and Accessibility; October 26–29, 2025; Denver, CO, USAccs: Human-centered computing Accessibility technologiesccs: Social and professional topics People with disabilitiesccs: Human-centered computing Interaction design process and methodsccs: Computing methodologies Mixed / augmented realityccs: Human-centered computing Ubiquitous and mobile devices
@@ 1. Introduction
Recent advances in extended reality (XR) technologies (Kasowski et al., 2023) and artificial intelligence (Vaswani et al., 2017; Yin et al., 2023; Wu et al., 2023) have opened new possibilities for vision-based assistive tools. In particular, head-mounted XR displays offer a promising form factor due to their visual, wearable, and hands-free nature (Kim and Choi, 2021). In this paper, we use the term “smart glasses” to specifically refer to head-mounted XR displays, such as the Apple Vision Pro (app, 2023). This excludes devices without XR capabilities, such as the Ray-Ban Meta glasses (Wikipedia contributors, 2024). Smart glasses have been widely explored for supporting people with ocular vision impairments (Li et al., 2022; Kasowski et al., 2023), including contrast enhancement, magnification, and navigational overlays (Zhao et al., 2015, 2016, 2020). While these technologies hold promise, their application to CVI remains largely unexplored.
@@
To address this, we conducted an eight month co-design study with two adults with CVI (who are also co-authors). Guided by the Double Diamond design framework (Banathy, 2013; Sanders and Stappers, 2008), we began with a two-week diary study to uncover their everyday challenges and management strategies. This was followed by individual ideation and scoping workshops identifying six representative challenges and potential design solutions. We then conducted ten iterative development workshops with our co-designers using the Apple Vision Pro to collaboratively prototype, test, and refine these solutions. Our specific contributions are:
@@ 2.1. Smart Glasses and XR for Visual Assistance
Kasowski et al. (Kasowski et al., 2023) further expanded this landscape by reviewing 76 XR-based studies targeting people with low vision. These studies explored a variety of visual augmentation strategies—ranging from contrast enhancement (Zhao et al., 2015, 2016, 2020) to spatial cues for navigation (Hommaru and Tanaka, 2020; Zhao et al., 2019)—and highlighted the growing potential of XR in augmentations for vision impairments. Importantly, Kasowski et al. called for more real-world validation and user-centred design approaches. However, their review did not include CVI, despite its growing recognition as a distinct form of brain-based visual impairment. This omission points to a broader gap: while smart glasses have been widely studied for ocular conditions, their application for CVI remains under-explored.
@@ 3. Methodology
This eight-month study followed the Double Diamond design framework (Design Council, 2005; Banathy, 2013), a structured, method-agnostic approach to design that emphasises user needs and iterative refinement. Prior work has shown that this approach aligns well with addressing the specific requirements from AT users (Chacon-Cifuentes et al., 2020; Priya and Upadhyay, 2021; Howard et al., 2024; Restianty et al., 2024; Zhang et al., 2024)-leading to better user satisfaction and more practical assistive tools (Howard et al., 2024; Chacon-Cifuentes et al., 2020). The Double Diamond comprises of four phases—Discover, Define, Develop, and Deliver—each representing alternating divergent and convergent stages of the design process. This paper is structured around these four phases, with each section detailing the corresponding methods.
Within this framework, we adopted a co-design approach, working with end-users as equal partners (Sanders and Stappers, 2008). As illustrated in Figure 1, combining co-design with the Double Diamond model helped us generate realistic insights (Discover), frame problems accurately (Define), ideate relevant solutions (Develop), and create usable outputs that are likely to be adopted (Deliver) (Morris et al., 2024). This collaborative process also helps ensure that those involved understand and feel ownership of the outcomes.
@@ 3.1. Co-designers
This study involved close collaboration with two adults with CVI. We intentionally chose to work with a small group to allow a deep, individualised understanding of each participant’s experiences and needs. This approach reflected that the study was an exploratory investigation aimed at surfacing key design challenges and assessing the feasibility of early prototypes. As noted in prior work (Morrison et al., 2021; Kamath et al., 2024; Schröder and Ghajargar, 2021), working closely with a small number of participants can yield rich, in-depth data—prioritising quality and nuance over quantity—which is especially valuable when designing highly personalised and context-aware assistive technologies.
@@ 3.2. Device Platform
Given the exploratory focus of this study—evaluating whether smart glasses with visual augmentations can meaningfully support people with CVI—we selected a VST HMD: the Apple Vision Pro (Wikipedia contributors, 2025b). Its ability to deliver high-fidelity, spatially coherent visual overlays was essential for our use cases and justified the trade-off in reduced direct visual clarity associated with VST technology.
@@
Finally, we referred to the device as ‘glasses’ throughout the study to help co-designers envision a future with a less obtrusive form factor, despite the current prototype being a bulkier headset.
Co-design was a collaborative process in which the two co-designers brought their lived experience, while the research team supported translation into workable technical solutions. While the research team selected the device platform (Apple Vision Pro), facilitated workshop sessions, and addressed feasibility constraints, the design decisions were the result of conversations and experimentation between the researchers and co-designers. Co-designers identified the most important challenges they faced, added to the initial set of possible solutions during the ideation and scoping workshops (for instance, suggesting the user tap on a section of the page to enhance when reading), and in subsequent workshops they played a central role in refining prototypes by critically evaluating them and suggesting modifications, such as subtly scaling the text when reading and where to position name tags and arrows. This feedback was enhanced by ensuring the co-designers were familiar with the device platform and its capabilities and limitations. This was the result of allowing the co-designers to spend considerable time using the device during the ideation workshops and in the subsequent workshops.
[…]
Refer to supplementary materials Figure2.csv for the table text.
@@ 4.2. Data Analysis
The collected diary entries were compiled and analysed. Dijana contributed 18 entries and Nicola 20, across their selected days. Three researchers independently categorised each entry by challenge type, which were then grouped into seven high-level categories, identified in (Gamage et al., 2024a). Any discrepancies in coding were discussed and resolved through consensus. Subsequently, two researchers conducted a deductive thematic analysis of the strategies co-designers used to manage challenges, identifying four primary strategy types:
@@ 4.3. Challenges
Beyond the nature of individual challenges, the contexts in which they occurred also varied significantly. Nicola’s entries were often situated in professional environments, including international travel and academic commitments, while Dijana’s challenges were primarily in domestic and caregiving contexts. This highlights the diverse situational demands placed on assistive technologies—requiring flexibility across domains of use.
@@ 4.4. Management Strategies
In contrast, Nicola was more inclined to seek help when needed, particularly during unfamiliar situations encountered while travelling internationally during the study. This is reflected in the following entries:
@@ 5. Ideation and Scoping Workshop
First, it supported the Define phase—part of the initial Problem diamond—by helping each co-designer articulate and refine the specific difficulties they face in daily life. Second, it initiated the Develop phase—the beginning of the Solution diamond—by exploring possible smart glasses–based solutions tailored to those difficulties.
@@ 5.1. Procedure
Each workshop consisted of a three session with each co-designers, facilitated by the primary researcher. The session was divided into four parts:
@@
Given our main goal was to explore how smart glasses can overcome difficulties faced the co-designers, we used these ideated possible solutions as a lens to identify which difficulties were perceived as most addressable through smart glasses.
We did this by asking each co-designer to rate the usefulness of all previously ideated solutions on a 3-point scale (1=not useful, 2=somewhat useful, 3=very useful). We then filtered the original list of difficulties—identified during the workshop phase—to include only those that had at least one associated solution rated as ‘very useful.’ This process prioritised difficulties that were both relevant to the co-designers and seen as addressable by smart glasses.
@@ 6.1. Procedure
At the beginning of each session, the first author presented the solution to the co-designers. This was done with an active discussion, supported by a set of guiding questions designed to prompt deeper engagement—for example, whether any visual elements should be modified or if the motion and behaviour of specific overlays felt intuitive. As the co-designers interacted with the system, the research team documented their feedback.
@@ 6.2. Design Exploration
Two design options were developed to support person recognition using both visual and auditory modalities. Techniques explored were inspired by prior work in XR that explores name tag placement on both objects (Azuma and Furmanski, 2003) and people (Wang et al., 2023; Stevens, 2018).
[…]
@@
We explored 14 design options across five solution areas aimed at guiding visual attention. Design options explored were inspired by other studies in the field, which have used similar methods for guiding visual search (Lu et al., 2013, 2012) and navigation (Koch et al., 2014).
@@
    Both emphasised the potential of the device to serve a proactive role in detecting physiological signs of stress and intervening before sensory overload occurs. Nicola noted a correlation between her CVI meltdowns and a drop in heart rate variability (HRV), a known biomarker of stress (Rowe et al., 1998; Immanuel et al., 2023; Kim et al., 2018, 2024). Dijana similarly reported physical exhaustion due to visual fatigue. These observations suggest that integrating biometric data from connected wearables, such as smartwatches, could enable real-time stress monitoring and early intervention.
@@ 6.3. Outlook and Optimism
This enthusiasm echoes broader findings, where OVI users reported improved independence and quality of life with smart glasses (Varshney et al., 2025). However, issues such as affordability and technical limitations remain barriers to broader adoption (World Health Organization, 2024).
@@ 7. Discussion
This section discusses the findings from this study and its implications for smart glass research for people with CVI. We aim to answer our primary research question: How can smart glasses alleviate current challenges faced by people with CVI when understanding and interacting with their immediate environment?
@@ 7.1. For which challenges are smart glasses useful?
For reading text, while prior work has explored smartphone-based text reading aids (Guo et al., 2016; Tan et al., 2024), few studies have examined in-situ text augmentation using AR (Zhao et al., 2015). Our work extends this by embedding both visual and auditory support directly into the physical reading environment—a combination both co-designers preferred.
@@
All techniques for guiding visual attention explored in our study involved manipulating virtual elements overlaid between the user’s eyes and the device platform’s pass-through video. This approach aligns with other studies in the field, which have used similar methods for guiding visual search (Lu et al., 2013, 2012) and navigation (Koch et al., 2014). However, prior research has investigated saliency modulation directly on the pass-through video to influence attention (Sutton et al., 2022, 2024; Veas et al., 2011; Mendez et al., 2010). Due to device platform limitations discussed in Section 3.2, we were unable to explore direct modification of the pass-through video. Future work should investigate the effectiveness of such approaches for supporting visual attention in people with CVI.
@@
For instance, the use of visual highlighting to locate items (C2- Locating) also proved useful for drawing attention to salient objects in busy visual scenes (C5-Sensory Overload). Visual markers used to guide or redirect attention (C5-Sensory Overload) may enhance situational awareness (C1-Unawareness), as suggested by prior work (Lu et al., 2013, 2012). Similarly, edge-based highlighting techniques could also benefit users with difficulties in subject and background separation (C7-Luminance and Contrast Sensitivity). Likewise, the use of 3D directional arrows (C2-Locating) were discussed by our co-designers as potentially suited to navigation (C6-Mobility), and the solutions we developed to support recognising people (C3-Recognising) was also envisioned for locating people, e.g., asking the glasses to “Find Wendy,” with the device displaying a name tag when Wendy was identified—bridging C3-Recognising and C2-Locating.
@@
Both co-designers in our study were diagnosed with simultanagnosia—a high-level visual difficulty characterised by an inability to perceive multiple objects simultaneously. This condition has been shown to impact all six of the selected difficulties explored in our study (Dalrymple et al., 2007; Baylis et al., 1994; Friedman-Hill et al., 1995; Thomas et al., 2009; Rizzo and Vecera, 2002; Karnath and Perenin, 2005). More recent findings also describe a dynamic constriction of visual attention under environmental complexity (St Clair Tracy et al., 2025), highlighting the evolving nature of how these conditions are understood.
@@
Because both co-designers in this study shared the same diagnosis, our design process and findings were naturally shaped around their specific perceptual experiences. As such, while our results suggest that certain visual strategies (e.g., strong colour contrast, spatial overlays) may benefit people with simultanagnosia, their applicability to other high-level visual impairments remains uncertain.
@@
Prior research highlights the role of neuroplasticity in reshaping visual processing through targeted interventions (Bennett et al., 2020; Delay et al., 2023; McDowell, 2023; Weden et al., 2022; Waddington and Hodgson, 2017; Ciman et al., 2013). For instance, prior studies on perceptual learning show that targeted exposure can improve face recognition by people with prosopagnosia (Byrne and Porter, 2025). While neuroplasticity alone cannot overcome CVI—a lifelong condition rooted in neurological differences—it can support adaptation by strengthening compensatory pathways and strategies over time.

### v3

@@ Abstract.
journalyear: 2025copyright: ccconference: The 27th International ACM SIGACCESS Conference on Computers and Accessibility; October 26–29, 2025; Denver, CO, USAbooktitle: The 27th International ACM SIGACCESS Conference on Computers and Accessibility (ASSETS ’25), October 26–29, 2025, Denver, CO, USAdoi: 10.1145/3663547.3746383isbn: 979-8-4007-0676-9/2025/10ccs: Human-centered computing Accessibility technologiesccs: Social and professional topics People with disabilitiesccs: Human-centered computing Interaction design process and methodsccs: Computing methodologies Mixed / augmented realityccs: Human-centered computing Ubiquitous and mobile devices
@@ 1. Introduction
Recent advances in extended reality (XR) technologies (Kasowski et al., 2023) and artificial intelligence (Vaswani et al., 2017; Yin et al., 2024; Wu et al., 2023) have opened new possibilities for vision-based assistive tools. In particular, head-mounted XR displays offer a promising form factor due to their visual, wearable, and hands-free nature (Kim and Choi, 2021). In this paper, we use the term “smart glasses” to specifically refer to head-mounted XR displays, such as the Apple Vision Pro (Apple Inc., 2023). This excludes devices without XR capabilities, such as the Ray-Ban Meta glasses (Wikipedia contributors, 2025d). Smart glasses have been widely explored for supporting people with ocular vision impairments (Li et al., 2022; Kasowski et al., 2023), including contrast enhancement, magnification, and navigational overlays (Zhao et al., 2015, 2016, 2020). While these technologies hold promise, their application to CVI remains largely unexplored.
@@
To address this, we conducted an eight month co-design study with two adults with CVI (who are also co-authors). Guided by the Double Diamond design framework (Banathy, 1996; Sanders and Stappers, 2008), we began with a two-week diary study to uncover their everyday challenges and management strategies. This was followed by individual ideation and scoping workshops identifying six representative challenges and potential design solutions. We then conducted ten iterative development workshops with our co-designers using the Apple Vision Pro to collaboratively prototype, test, and refine these solutions. Our specific contributions are:
@@ 2.1. Smart Glasses and XR for Visual Assistance
Kasowski et al. (Kasowski et al., 2023) further expanded this landscape by reviewing 76 XR-based studies targeting people with low vision. These studies explored a variety of visual augmentation strategies—ranging from contrast enhancement (Zhao et al., 2015, 2016, 2020) to spatial cues for navigation (Hommaru and Tanaka, 2020; Zhao et al., 2019)—and highlighted the growing potential of XR in augmentations for vision impairments. Importantly, Kasowski et al. called for more real-world validation and user-centred design approaches. However, their review did not include CVI, despite its growing recognition as a brain-based visual impairment. This omission points to a broader gap: while smart glasses have been widely studied for ocular conditions, their application for CVI remains under-explored.
@@ 3. Methodology
This eight-month study followed the Double Diamond design framework (Design Council, 2005; Banathy, 1996), a structured, method-agnostic approach to design that emphasises user needs and iterative refinement (Priya and Upadhyay, 2021; Howard et al., 2024; Restianty et al., 2024; Zhang et al., 2024). This paper is structured around the four phases of the Double Diamond—Discover, Define, Develop, and Deliver—which represent alternating divergent and convergent stages of the design process, with each section detailing the methods used. Within this framework, we adopted a co-design approach (Sanders and Stappers, 2008; Morris et al., 2024) , engaging end-users as equal partners to support shared understanding and ownership of the outcomes. Figure 1 summarises the methodology.
@@ 3.1. Co-designers
This study involved close collaboration with two adults with CVI. We intentionally chose to work with a small group to allow for rich, in-depth and individualised data—prioritising quality and nuance over quantity—which is especially valuable when designing highly personalised and context-aware assistive technologies (Morrison et al., 2021; Kamath et al., 2024; Schröder and Ghajargar, 2021). This approach reflected that the study was an exploratory investigation aimed at surfacing key design challenges and assessing the feasibility of early prototypes.
@@
Co-design in this project was a collaborative process in which the two co-designers brought their lived experience, while the research team supported translation into workable technical solutions. While the research team selected the device platform, facilitated workshop sessions, and addressed feasibility constraints, the design decisions were the result of conversations and experimentation between the researchers and co-designers. Co-designers identified the most important challenges they faced (Section 4) and added to the initial set of possible solutions during the ideation and scoping workshops (Section 5). For instance, the co-designers suggested the user tap on a section of the page to enhance when reading. In subsequent workshops (Section 6), they played a central role in refining prototypes by critically evaluating them and suggesting modifications, such as subtly scaling the text when reading and where to position name tags and arrows. This feedback was enhanced by ensuring the co-designers were familiar with the device platform and its capabilities and limitations. This was the result of allowing the co-designers to spend considerable time using the device during the ideation workshops and in the subsequent workshops.
@@ 3.2. Device Platform
Given the exploratory focus of this study we selected a VST HMD: the Apple Vision Pro (Wikipedia contributors, 2025b). Its ability to deliver high-fidelity, spatially coherent visual overlays was essential for our use cases and justified the trade-off in reduced direct visual clarity associated with VST technology.
@@
We referred to the device as ‘glasses’ throughout the study to help co-designers envision a future with a less obtrusive form factor, despite the current prototype being a bulkier headset.
@@ 4.2. Data Analysis
The collected diary entries were compiled and analysed. Dijana contributed 18 entries and Nicola 20. Three researchers independently categorised each entry by challenge type, which were then grouped into seven high-level categories, identified in (Gamage et al., 2024a). Any discrepancies in coding were discussed and resolved through consensus. Subsequently, two researchers conducted a deductive thematic analysis of the strategies co-designers used to manage challenges, identifying four primary strategy types:
@@
[…]
Refer to supplementary materials Figure2.csv for the table text.
@@ 4.3. Challenges
Beyond the nature of individual challenges, the contexts in which they occurred also varied significantly. Nicola’s entries were often situated in professional environments, including air travel and academic commitments, while Dijana’s challenges were primarily in domestic and caregiving contexts. This highlights the diverse situational demands placed on assistive technologies—requiring flexibility across domains of use.
@@ 4.4. Management Strategies
In contrast, Nicola was more inclined to seek help when needed, particularly during unfamiliar situations encountered while travelling during the study. This is reflected in the following entries:
@@ 5. Ideation and Scoping Workshop
First, it supported the Define phase by helping each co-designer articulate and refine the specific difficulties they face in daily life. Second, it initiated the Develop phase by exploring possible smart glasses–based solutions tailored to those difficulties.
@@ 5.1. Procedure
Each workshop consisted three sessions with each co-designers, facilitated by the primary researcher. The session was divided into four parts:
@@
To identify which solutions were of greatest potential benefit to our co-designers, they rated the usefulness of all previously ideated solutions on a 3-point scale (1=not useful, 2=somewhat useful, 3=very useful). We then filtered the original list of difficulties—identified during the workshop phase—to include only those that had at least one associated solution rated as ‘very useful’. This process prioritised difficulties that were both relevant to the co-designers and seen as addressable by smart glasses.
@@ 6.1. Procedure
At the beginning of each session, the first author presented the solution to the co-designers. This was done with an active discussion, supported by a set of guiding questions designed to prompt deeper engagement—for example, whether any visual elements should be modified or if the motion and behaviour of specific overlays felt intuitive. Feedback and discussions were recorded while the co-designers interacted with the system.
@@ 6.2. Design Exploration
Two design options were developed to support person recognition using both visual and auditory modalities. Techniques explored were inspired by prior work in XR that explores name tag placement on both objects (Azuma and Furmanski, 2003) and people (Wang et al., 2023; Stevens, 2019).
@@
We explored 14 design options across five solution areas aimed at guiding visual attention. Design options explored were inspired by other studies in the field, which have used similar methods for guiding visual search (Lu et al., 2012) and navigation (Koch et al., 2014).
@@
Figure 10. Summary of findings from the development workshops for the six identified difficulties.
[…]
Image showing summary of findings from the development workshops for the six identified difficulties, each includes a title, image, and two to three findings lablled as liked and disliked. Locating Items in a Cluttered Environment Image: a cluttered surface with paper, bottles, and containers. One cup is highlighted in blue. liked: Highlight or outline the object while dimming the environment liked: Large arrow when the object is out of view disliked: Highlighting general area or head movement
Reading Text image: a hand points to a section of a printed document. One line is bolded where the index finger is. liked: Text to speech by tapping on a section of the page liked: Enhance a line of text using the finger disliked: Hand-tracking precision due to optic ataxia
Face-to-face Conversations image: two people face each other; the person being spoken to has a blue box around their upper body and head. liked: Fixed box to the upper body while dimming the environment liked: Switch to person who is talking automatically disliked: Dynamic speaker silhouette is visually distracting
Recognising People image: a person faces forward with a virtual blue name tag labelled “John” floating under their chin. liked: Virtual name tag closer to the user’s face liked: Fade away once glanced at the name tag disliked: Only detect pre-selected people to avoid visual clutter
Visual Attention image: a cluttered tabletop with documents, cups, and a bottle. Blue dots indicate markers to guide visual attention. liked: Education tool to help teach how to guide visual attention disliked: Not an effective real-time assistive tool for guiding visual attention, as it causes discomfort and sensory overload in everyday use
Stressed due to Sensory Overload image: a person wears a headset showing a calming virtual scene with mountains and clouds. liked: Help manage stress by offering visual disengagement through immersive environments liked: Integration with biometric signals like HRV, eye movement can enable proactive support
@@
    Both emphasised the potential of the device to serve a proactive role in detecting physiological signs of stress and intervening before sensory overload occurs. Nicola noted a correlation between her CVI meltdowns and a drop in heart rate variability (HRV), a known biomarker of stress (Rowe et al., 1998; Immanuel et al., 2023; Hye-Geum et al., 2018; Kim et al., 2024). Dijana similarly reported physical exhaustion due to visual fatigue. These observations suggest that integrating biometric data from connected wearables, such as smartwatches, could enable real-time stress monitoring and early intervention.
@@
  • [Uncaptioned image]
    Since full visual block is only feasible when stationary, both co-designers preferred guided sensory reduction with adjustable modes—Minimal, Medium, and Extreme—that could adapt to different contexts. For Minimal, they preferred options like white noise or darkening the environment with ambient environment sound. Medium involved fully darkening the environment or using the Mindfulness app. Extreme combined immersive visual environments with noise cancellation. They expressed interest in customising each mode as they became more familiar with the glasses.
@@ 6.3. Outlook and Optimism
This enthusiasm echoes broader findings, where OVI users reported improved independence and quality of life with smart glasses (Varshney et al., 2024). However, issues such as affordability and technical limitations remain barriers to broader adoption (World Health Organization, 2024).
@@ 7. Discussion
This section discusses the findings from the study (summarised in Figure 10) and explores their implications for smart glass research for people with CVI. We aim to answer our primary research question: How can smart glasses alleviate current challenges faced by people with CVI when understanding and interacting with their immediate environment?
@@ 7.1. For which challenges are smart glasses useful?
For reading text, while prior work has explored smartphone-based text reading aids (Guo et al., 2016; Tan et al., 2022), few studies have examined in-situ text augmentation using AR (Zhao et al., 2015). Our work extends this by embedding both visual and auditory support directly into the physical reading environment—a combination both co-designers preferred.
@@
All techniques for guiding visual attention explored in our study involved manipulating virtual elements overlaid between the user’s eyes and the device platform’s pass-through video. This approach aligns with other studies in the field, which have used similar methods for guiding visual search (Lu et al., 2012) and navigation (Koch et al., 2014). However, prior research has investigated saliency modulation directly on the pass-through video to influence attention (Sutton et al., 2022, 2024; Veas et al., 2011; Mendez et al., 2010). Due to device platform limitations discussed in Section 3.2, we were unable to explore direct modification of the pass-through video. Future work should investigate the effectiveness of such approaches for supporting visual attention in people with CVI.
@@
For instance, the use of visual highlighting to locate items (C2- Locating) also proved useful for drawing attention to salient objects in busy visual scenes (C5-Sensory Overload). Visual markers used to guide or redirect attention (C5-Sensory Overload) may enhance situational awareness (C1-Unawareness), as suggested by prior work (Lu et al., 2012). Similarly, edge-based highlighting techniques could also benefit users with difficulties in subject and background separation (C7-Luminance and Contrast Sensitivity). Likewise, the use of 3D directional arrows (C2-Locating) were discussed by our co-designers as potentially suited to navigation (C6-Mobility), and the solutions we developed to support recognising people (C3-Recognising) was also envisioned for locating people, e.g., asking the glasses to “Find Wendy,” with the device displaying a name tag when Wendy was identified—bridging C3-Recognising and C2-Locating.
@@
Both co-designers in our study were diagnosed with simultanagnosia—a high-level visual difficulty characterised by an inability to perceive multiple objects simultaneously. This condition has been shown to impact all six of the selected difficulties explored in our study (Dalrymple et al., 2007; Baylis et al., 1994; Friedman-Hill et al., 1995; Thomas et al., 2009; Rizzo and Vecera, 2002; Karnath and Perenin, 2005). More recent findings also describe a dynamic constriction of visual attention under environmental complexity (St Clair Tracy et al., 2024), highlighting the evolving nature of how these conditions are understood.
@@
Because both co-designers in this study shared a similar diagnosis, our design process and findings were naturally shaped around their specific perceptual experiences. As such, while our results suggest that certain visual strategies (e.g., strong colour contrast, spatial overlays) may benefit people with simultanagnosia, their applicability to other high-level visual impairments remains uncertain.
@@
Prior research highlights the role of neuroplasticity in reshaping visual processing through targeted interventions (Bennett et al., 2020; Delay et al., 2022; McDowell, 2023; Weden et al., 2022; Waddington and Hodgson, 2017; Ciman et al., 2013). For instance, prior studies on perceptual learning show that targeted exposure can improve face recognition by people with prosopagnosia (Byrne and Porter, 2025). While neuroplasticity alone cannot overcome CVI—a lifelong condition rooted in neurological differences—it can support adaptation by strengthening compensatory pathways and strategies over time.